OpenStudy (anonymous):

BEST ANSWER GETS THE MEDAL!!!!!!!!!!!!!!!!!! For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

OpenStudy (anonymous):

@hartnn @.Sam. @amistre64

OpenStudy (anonymous):

@ivettef365 can you help me?

OpenStudy (anonymous):

two beers and one shot cost $10 three beers and two shots cost $25 if \(x\) is the number of beers and \(y\) is the number of shots how much does each cost?

OpenStudy (anonymous):

i should say "if \(x\) is the price of a beer, and \(y\) is the price of a shot, how much does each cost?"

OpenStudy (anonymous):

you can solve by rewriting \(2x+y=10\) as \(y=10-2x\) and then substitute in to the second equation to get \[3x+4(10-2x)=25\]

OpenStudy (anonymous):

if you think "beer and shots" will annoy your teacher, then you can use "hotdogs and hamburgers"

OpenStudy (anonymous):

ok thanks your a life saver!!!!!!! but i still have to slove the problem right